OpenStudy (anonymous):

Laplace transform for f(t)=tcosh(3t)?

OpenStudy (anonymous):

I looked at the table and it says (-1)^n*F^(n) of (s)

OpenStudy (anonymous):

the laplace transform of cosh(3t) is s/(s^2-9) so I thought it was that multiplied by -1

OpenStudy (anonymous):

but why is the answer taking the derivative of it

OpenStudy (jamesj):

No, because L[fg] does not equal L[f]L[g] in general. (That's the formula for convolution.) I'd write cosh(3t) = 1/2 (e^3t + e^-3t). Then find the Laplace transform of te^at and then use linear superposition.

OpenStudy (anonymous):

I don't understand.. is 1/2 (e^3t + e^-3t) that just another way to write cosh(3t)?

OpenStudy (jamesj):

yes, that's actually the definition of cosh

OpenStudy (jamesj):

and sinh(at) = 1/2 (e^at - e^-at)

OpenStudy (anonymous):

Oh I guess that works too, but could you tell me why they took the different in this solution?

OpenStudy (anonymous):

derivative**

OpenStudy (jamesj):

Yes, it's because \[ L[ t g(t) ] = - G'(s) \] where \( L[g(t)] = G(s) \) That identity is true because \[ \frac{d \ }{ds} G(s) = \frac{d \ }{ds} \int_0^\infty g(t) e^{-st} dt \] \[ = \int_0^\infty \frac{\partial \ }{\partial s} ( g(t) e^{-st}) \ dt \] \[ = \int_0^\infty g(t) \frac{\partial \ }{\partial s} e^{-st} \ dt \] \[ = \int_0^\infty g(t) . (-t) e^{-st} \ dt \] \[ = -\int_0^\infty t g(t) . e^{-st} \ dt \] \[ = -L[tg(t)] \]

OpenStudy (anonymous):

Oh I see it make sense now thanks! I actually have another question but I'll post it in a new board. Thanks agian